# Intersection and Union Hierarchies of Deterministic

### v2

@@ 4 Proofs of Three Separation Claims
The first pumping lemma for $\mathrm{DCFL}[2^{k-1}-1]$ provides a constant $c>0$ and a distinct pair $s,s^{\prime}\in I_{k}$. Let $s=s_{2}s_{4}\cdots s_{k-2}s_{k}s_{k-1}\cdots s_{3}s_{1}$ and $s^{\prime}=s^{\prime}_{2}s^{\prime}_{4}\cdots s^{\prime}_{k-2}s^{\prime}_{k}s^{\prime}_{k-1}\cdots s^{\prime}_{3}s^{\prime}_{1}$. We take the minimum even number $j$ satisfying $s_{j}\neq s^{\prime}_{j}$ and then set $n_{j}$ to be sufficiently larger than $c$. Let us consider two corresponding strings $\alpha_{s}$ and $\alpha_{s^{\prime}}$. There are 6 cases to consider separately. Our proof strategy is that, in each such case, by choosing $\hat{x}$, $\hat{y}$, and $\hat{z}$ properly, we can define $x^{\prime}=\alpha_{s}^{(j)}\hat{x}$, $y=\hat{y}\beta_{s}^{(j)}$, and $z=\hat{z}\beta_{s}^{(j)}$ to form $\alpha_{s}=x^{\prime}y$ and $\alpha_{s^{\prime}}=x^{\prime}z$. We then apply the conditions (1)–(2) of the pumping lemma to draw a contradiction.
(i) If $(s_{j},s^{\prime}_{j})=(1,2)$, we set $\hat{x}=a^{n_{j}}b^{n_{j}}$, $\hat{y}=c^{n_{j}}$, and $\hat{z}=bc^{n_{j}}$. (ii) If $(s_{j},s^{\prime}_{j})=(3,4)$, we set $\hat{x}=a^{n_{j}}b^{n_{j}-1}c^{n_{j}-1}$, $\hat{y}=\varepsilon$, and $\hat{z}=c$. (iii) If $(s_{j},s^{\prime}_{j})=(1,3)$, then we set $\hat{x}=a^{n_{j}}$, $\hat{y}=b^{n_{j}}c^{n_{j}-1}$, and $\hat{z}=b^{n_{j}-1}c^{n_{j}-1}$. (iv) If $(s_{j},s^{\prime}_{j})=(1,4)$, then we set $\hat{x}=a^{n_{j}}b^{n_{j}-1}$, $\hat{y}=bc^{n_{j}-1}$, and $\hat{z}=c^{n_{j}}$. (v) If $(s_{j},s^{\prime}_{j})=(2,3)$, then we set $\hat{x}=a^{n_{j}}$, $\hat{y}=b^{n_{j}+1}c^{n_{j}}$, and $\hat{z}=b^{n_{j}-1}c^{n_{j}-1}$. (vi) If $(s_{j},s^{\prime}_{j})=(2,4)$, then we set $\hat{x}=a^{n_{j}}b^{n_{j}-1}$, $\hat{y}=b^{2}c^{n_{j}}$, and $\hat{z}=c^{n_{j}}$.
@@
Let us consider the case (i). Since $\hat{y}$ contains only one $b$, the matching symbol in $\hat{x}$ is $b$ in $b^{n_{j}-1}$ to form an iterative pair of $x^{\prime}y$ for $L_{k}$. However, since $\hat{z}$ has no $b$, there is no iterative pair of $x^{\prime}z$ for $L_{k}$, a contradiction. Next, we consider the case (iii). Consider an iterative pair $(a,b)$ of $x^{\prime}y$ for $L_{k}$, where $a$ is in $\hat{x}$ and $b$ is in $\hat{y}$. However, we cannot pump $a$ in $x^{\prime}z$, a contradiction. For the other cases, we can use a similar argument to draw a contradiction. This completes the proof of the proposition. ∎
@@ 5.2 State-Stack Pairs and Mutual Correlations
    Let $t_{1},t_{2},t_{3},t_{4}\in\mathbb{N}$ with $1\leq t_{1}<t_{2}<t_{3}<t_{4}\leq|w|+1$. Let $w=x_{1}x_{2}x_{3}x_{4}x_{5}$ so that each $t_{i}$ is an $(x_{i},x_{i+1})$-boundary for each index $i\in[4]$. If the pair $([t_{1},t_{2}],[t_{3},t_{4}])$ is good, then $(x_{2},x_{4})$ is an iterative pair of $w$ for $L(M)$.
@@
We assume that the condition (1) fails. Our goal is then to verify that the condition (2) of the lemma indeed holds. There are four specific cases to deal with. Hereafter, we intend to discuss them separately. Remember that, since $M$’s tape head moves in only one direction, any state-stack pair at each boundary in the $x$-region does not depend on the choice of $y$ and $z$.
@@
Case 3: For convenience, we set R=(|x|c,|x|)R=(|x|-c,|x|). We first consider the simple case where there is no pop operation in the $R$-region. We can take a number $m\geq 1$ and a series of $m$ boundaries $s_{1},s_{2},\ldots,s_{m}$ in the $R$-region with $1\leq s_{1}<s_{2}<\cdots<s_{m}$ such that, for appropriately chosen elements $q\in Q$, $\gamma\in(\Gamma^{(-)})^{*}Z_{0}$, and $\alpha_{1},\ldots,\alpha_{m-1}\in(\Gamma^{(-)})^{*}$, there exist state-stack pairs of the form $(q,\gamma),(q,\alpha_{1}\gamma),\ldots,(q,\alpha_{[m-1,1]}\gamma)$ respectively at the boundaries $s_{1},s_{2}\ldots,s_{m}$, where $\alpha_{[i,1]}$ is the abbreviation of $\alpha_{i}\alpha_{i-1}\cdots\alpha_{1}$ introduced in Section 5.2. Since no pop operation is performed, the boundary blocks $s_{1},s_{2}],[s_{2},s_{3}],\ldots,[s_{m-1},s_{m}]$ are all convex. By maximizing the value of $m$, we hereafter assume that $m$ takes this maximum value. Notice that $R$-region contains more than $|Q|^{3}$ boundaries because of $|Q|^{3}<c$. If, for each $q\in Q$, the corresponding value $m_{q}$ is at most $|Q|^{2}$, then the size $c$ of the $R$-region must be at most $|Q|^{3}$. This is a contradiction, and thus we conclude that $m>|Q|^{2}$.
@@
Next, we choose two series $\{t_{i}\}_{i\in[m]}$ and $\{r_{i}\}_{i\in[m]}$ of boundaries in the $y$-region and the $z$-region, respectively, for which (a) $s_{m}\leq t_{1}<t_{2}<\cdots<t_{m}$ and $s_{m}\leq r_{1}<r_{2}<\cdots<r_{m}$ and (b) for any index $i\in[m-1]$, $s_{i},s_{i+1}]$ is mutually correlated to $t_{m-i},t_{m-i+1}]$ in the $y$-region and also to $r_{m-i},r_{m-i+1}]$ in the $z$-region. This is possible because $M$ is $\varepsilon$-enhanced and in an ideal shape. Notice that the boundary blocks $t_{1},t_{2}],\ldots,[t_{m-1},t_{m}]$ and $r_{1},r_{2}],\ldots,[r_{m-1},r_{m}]$ are all pseudo-convex. Assume that, at the boundaries $t_{1},t_{2},\ldots,t_{m}$, the associated state-stack pairs are respectively of the form $(p_{1},\alpha_{[m-1,1]}\gamma),(p_{2},\alpha_{[m-2,1]}\gamma),\ldots,(p_{m-1},\alpha_{1}\gamma),(p_{m},\gamma)$ for appropriate inner states $p_{1},p_{2},\ldots,p_{m}\in Q$. Similarly, assume that $(e_{1},\alpha_{[m-1,1]}\gamma),(e_{2},\alpha_{[m-2,1]}\gamma),\ldots,(e_{m-1},\alpha_{1}\gamma),(e_{m},\gamma)$ are state-stack pairs at the boundaries $r_{1},r_{2},\ldots,r_{m}$, respectively.
@@
By Lemma 5.3(2), since $m>|Q|^{2}$, there is a subset $C$ of $[m]$ of size more than $|Q|$ such that, for any index pair $i,j\in C$, the inner states at the boundaries $t_{i}$ and $t_{j}$ coincide; namely, $p_{i}=p_{j}$. From this fact, it follows that there is a special pair $j_{1},j_{2}\in C$ with $j_{1}<j_{2}$ for which the inner states at the boundaries $r_{j_{1}}$ and $r_{j_{2}}$ coincide; that is, $e_{j_{1}}=e_{j_{2}}$. We fix such a pair $(j_{1},j_{2})$. We then factorize the strings $x$, $y$, and $\hat{y}$ as $x=x_{1}x_{2}x_{3}$, $y=y_{1}y_{2}y_{3}$, and $z=z_{1}z_{2}z_{3}$ so that $s_{j_{1}}$ is the $(x_{1},x_{2})$-boundary, $s_{j_{2}}$ is the $(x_{2},x_{3})$-boundary, $t_{j_{1}}$ is the $(y_{1},y_{2})$-boundary, $t_{j_{2}}$ is the $(y_{2},y_{3})$-boundary, $r_{j_{1}}$ is the $(z_{2},z_{3})$-boundary, and $r_{j_{2}}$ is the $(z_{2},z_{3})$-boundary. It then follows that $(x_{2},y_{2})$ and $(x_{2},z_{2})$ are respectively iterative pairs of $xy$ and $xz$ for $L$. Therefore, the condition (2) follows.
@@
(a) We first consider the case where $s,s^{\prime}]$ is a single turn. We focus on a hill, say, $t_{1},t_{2}]$ that is contained in the turn and claim that this hill has height at most $|Q|^{2}$. This is shown as follows. If the hill has height more than $|Q|^{2}$, then there exists an inner state $q\in Q$ that appears in at least $|Q|+1$ state-stack pairs in the left area of the hill. Let $(q,\gamma),(q,\alpha_{1}\gamma),\ldots,(q,\alpha_{[|Q|,1]}\gamma)$ denote those state-stack pairs for appropriately chosen strings $\gamma,\alpha_{1},\ldots,\alpha_{|Q|+1}$. Associated with them, we also choose state-stack pairs $(p_{1},\alpha_{[|Q|,1]}\gamma),\ldots,(p_{|Q|},\alpha_{1}\gamma),(p_{|Q|+1},\gamma)$ in the right area of the hill for certain inner states $p_{1},\ldots,p_{|Q|+1}$. Lemma 5.2(1) then guarantees the existence of a distinct pair $i,j\in[|Q|+1]$ satisfying $p_{i}=p_{j}$. This leads to Case 2, a contradiction.
@@
(3) We estimate the total number of the turns whose bottom heights are a fixed number $h$. Given such a number $h$, we define $bh(h)$ to be the set of all indices $i\in[m-1]$ for which the bottom height of $s_{i},s_{i+1}]$ is exactly $h$. Note that $|bh(0)|\leq|Q|$ follows by (2b). We first assume that $|bh(h+1)|>|\bigcup_{i\in[0,h]_{\mathbb{Z}}}bh(i)|\cdot|Q|$ holds for a certain number $h\in\mathbb{N}$. This implies the existence of more than $|Q|$ turns in $bh(h+1)$, any pair of which contains no turn in $\bigcup_{i\in[0,h]_{\mathbb{Z}}}bh(i)$. Consider $|Q|+1$ state-stack pairs at the boundaries whose height matches the bottom heights of those specific turns. By Lemma 5.3(1), at least two of those state-stack pairs coincide, a contradiction. Hence, we conclude that, for any value $h\in\mathbb{N}$, $|bh(h+1)|\leq|\bigcup_{i\in[0,h]_{\mathbb{Z}}}bh(i)|\cdot|Q|$. We then assert that $|bh(h)|\leq 2^{h-1}|Q|^{h+1}$ for any $h\in\mathbb{N}^{+}$. Since $|bh(i)|\leq 2^{i-1}|Q|^{i+1}$ holds for any $i\in\mathbb{N}^{+}$, by induction hypothesis, it follows that $|bh(h)|\leq|Q|\cdot|\bigcup_{i\in[0,h-1]_{\mathbb{Z}}}bh(i)|\leq|Q|(|Q|+\sum_{i=1}^{h-1}2^{i-1}|Q|^{i+1})\leq(1+\sum_{i=1}^{h-1}2^{i-1})|Q|^{h+1}=2^{h-1}|Q|^{h+1}$.
@@
Basis Case of $d=1$: We begin our proof for the basis case of $d=1$. This comes from Lemma 3.2 by taking a finite-state 1dpda $M$. This completes the proof of the basis case of $d=1$.

### v3

@@ 4 Proofs of Three Separation Claims
The second pumping lemma for $\mathrm{DCFL}[2^{\lfloor k/2\rfloor}-1]$ provides a constant $c>0$ and a distinct pair $s,s^{\prime}\in A_{vec}$. Let $s=s_{2}s_{4}\cdots s_{k-2}s_{k}s_{k-1}\cdots s_{3}s_{1}$ and $s^{\prime}=s^{\prime}_{2}s^{\prime}_{4}\cdots s^{\prime}_{k-2}s^{\prime}_{k}s^{\prime}_{k-1}\cdots s^{\prime}_{3}s^{\prime}_{1}$. We take the minimum even number $j$ satisfying $s_{j}\neq s^{\prime}_{j}$. Without loss of generality, we assume that $s_{j}<s^{\prime}_{j}$. Let us consider two corresponding strings $\alpha_{s}$ and $\alpha_{s^{\prime}}$. For any even number $j\in[k-2]$, we define $\alpha_{s}^{(j)}=w_{2}^{(s_{2})}\#w_{4}^{(s_{4})}\#\cdots\#w_{j}^{(s_{j})}\#$ and $\beta_{s}^{(j)}=w_{j+2}^{(s_{j+2})}\#\cdots\#w_{3}^{(s_{3})}\#w_{1}^{(s_{1})}$. When $j=k$, we instead set $\beta_{s}^{(j)}=w_{j-1}^{(s_{j-1})}\#\cdots\#w_{3}^{(s_{3})}\#w_{1}^{(s_{1})}$. Notice that $\alpha_{s}=\alpha_{s}^{(j)}\beta_{s}^{(j)}$ holds.
Our proof strategy is that, by choosing $\hat{x}$, $\hat{y}$, and $\hat{z}$ properly, we define $x^{\prime}=\alpha_{s}^{(j-2)}\#\hat{x}$, $y=\hat{y}\#\beta_{s}^{(j)}$, and $z=\hat{z}\#\beta_{s}^{(j)}$ to form $\alpha_{s}=x^{\prime}y$ and $\alpha_{s^{\prime}}=x^{\prime}z$. We then apply the conditions (1)–(5) of the second pumping lemma to draw a contradiction. [^7] For this purpose, we actually need to slightly modify the original lemma so that the conditions (3)–(4) always hold among all degenerate iterative pairs satisfying the conditions without any length requirement on $x_{2}$ and $x^{\prime}_{2}$.
There are four cases to consider separately. (i) If $(s_{j},s^{\prime}_{j})=(1,3)$, then we set $\hat{x}=a^{n_{j}}b^{n_{j}-1}$, $\hat{y}=bc^{n_{j}-1}$, and $\hat{z}=c^{n_{j}-1}$. (ii) If $(s_{j},s^{\prime}_{j})=(1,4)$, then we set $\hat{x}=a^{n_{j}}b^{n_{j}-1}$, $\hat{y}=bc^{n_{j}-1}$, and $\hat{z}=c^{n_{j}}$. (iii) If $(s_{j},s^{\prime}_{j})=(2,3)$, then we set $\hat{x}=a^{n_{j}}b^{n_{j}-1}$, $\hat{y}=b^{2}c^{n_{j}}$, and $\hat{z}=c^{n_{j}-1}$. (iv) If $(s_{j},s^{\prime}_{j})=(2,4)$, then we set $\hat{x}=a^{n_{j}}b^{n_{j}-1}$, $\hat{y}=b^{2}c^{n_{j}}$, and $\hat{z}=c^{n_{j}}$.
@@
Finally, let us discuss the validity of the case (i). Consider any degenerate iterative pair $(x^{\prime}_{2},z_{2})$ of $x^{\prime}z$ for $L_{k}$, where $x^{\prime}_{2}$ is in $\hat{x}$ with $|x^{\prime}_{2}|\leq c$ and $z_{2}$ is in $\hat{z}$. However, there is no degenerate iterative pair for $x^{\prime}y$ in $\hat{x}\hat{y}$, a contradiction with the condition (4). For the other cases, we can use a similar argument to draw a contradiction.
This completes the proof of the proposition. ∎
@@ 5.2 State-Stack Pairs and Mutual Correlations
    Let $t_{1},t_{2},t_{3},t_{4}\in\mathbb{N}$ with $1\leq t_{1}<t_{2}<t_{3}<t_{4}\leq|w|+1$. Let $w=x_{1}x_{2}x_{3}x_{4}x_{5}$ so that each $t_{i}$ is the $(x_{i},x_{i+1})$-boundary for each index $i\in[4]$. If the pair $([t_{1},t_{2}],[t_{3},t_{4}])$ is good, then $(x_{2},x_{4})$ is an iterative pair of $w$ for $L(M)$.
@@
We assume that the condition (1) of the lemma fails. Our goal is then to verify that the condition (2) of the lemma indeed holds. There are four specific cases to deal with. Hereafter, we intend to discuss them separately. Remember that, since $M$’s tape head moves in only one direction, any state-stack pair at each boundary in the $x$-region does not depend on the choice of $y$ and $z$.
@@
Case 3: For convenience, we set R=(|x|c,|x|)R=(|x|-c,|x|). We first consider the simple case where there is no pop operation in the $R$-region. We can take a number $m\geq 1$ and a series of $m$ boundaries $s_{1},s_{2},\ldots,s_{m}$ in the $R$-region with $1\leq s_{1}<s_{2}<\cdots<s_{m}$ such that, for appropriately chosen elements $q\in Q$, $\gamma\in(\Gamma^{(-)})^{*}Z_{0}$, and $\alpha_{1},\ldots,\alpha_{m-1}\in(\Gamma^{(-)})^{*}$, there exist state-stack pairs of the form $(q,\gamma),(q,\alpha_{1}\gamma),\ldots,(q,\alpha_{[m-1,1]}\gamma)$ respectively at the boundaries $s_{1},s_{2}\ldots,s_{m}$, where $\alpha_{[i,1]}$ is the abbreviation of $\alpha_{i}\alpha_{i-1}\cdots\alpha_{1}$ introduced in Section 5.2. Since no pop operation is performed, the boundary blocks $s_{1},s_{2}],[s_{2},s_{3}],\ldots,[s_{m-1},s_{m}]$ are all convex. By maximizing the value of $m$, we hereafter assume that $m$ takes this maximum value. Notice that $R$-region contains more than $|Q|^{3}$ boundaries because of $|Q|^{3}<c$. If, for each $q\in Q$, the corresponding maximum value $m_{q}$ is at most $|Q|^{2}$, then the size $c$ of the $R$-region must be at most $|Q|^{3}$. This is a contradiction, and thus we conclude that $m>|Q|^{2}$.
@@
Next, we choose two series $\{t_{i}\}_{i\in[m]}$ and $\{r_{i}\}_{i\in[m]}$ of boundaries in the $y$-region and the $z$-region, respectively, for which (a) $s_{m}\leq t_{1}<t_{2}<\cdots<t_{m}$ and $s_{m}\leq r_{1}<r_{2}<\cdots<r_{m}$ and (b) for any index $i\in[m-1]$, $s_{i},s_{i+1}]$ is mutually correlated to $t_{m-i},t_{m-i+1}]$ in the $y$-region and also to $r_{m-i},r_{m-i+1}]$ in the $z$-region. This is possible because $M$ is $\varepsilon$-enhanced and in an ideal shape. Notice that the boundary blocks $t_{1},t_{2}],\ldots,[t_{m-1},t_{m}]$ and $r_{1},r_{2}],\ldots,[r_{m-1},r_{m}]$ are all pseudo-convex. Assume that, at the boundaries $t_{1},t_{2},\ldots,t_{m}$, the associated state-stack pairs are respectively of the form $(p_{1},\alpha_{[m-1,1]}\gamma),(p_{2},\alpha_{[m-2,1]}\gamma),\ldots,(p_{m-1},\alpha_{1}\gamma),(p_{m},\gamma)$ for appropriate inner states $p_{1},p_{2},\ldots,p_{m}\in Q$. Similarly, assume that $(e_{1},\alpha_{[m-1,1]}\gamma),(e_{2},\alpha_{[m-2,1]}\gamma),\ldots,(e_{m-1},\alpha_{1}\gamma),(e_{m},\gamma)$ are respectively state-stack pairs at the boundaries $r_{1},r_{2},\ldots,r_{m}$.
@@
By Lemma 5.3(2), since $m>|Q|^{2}$, there is a subset $C$ of $[m]$ of size more than $|Q|$ such that, for any index pair $i,j\in C$, the inner states at the boundaries $t_{i}$ and $t_{j}$ coincide; namely, $p_{i}=p_{j}$. From this fact, it follows that there is a special pair $j_{1},j_{2}\in C$ with $j_{1}<j_{2}$ for which the inner states at the boundaries $r_{j_{1}}$ and $r_{j_{2}}$ coincide; that is, $e_{j_{1}}=e_{j_{2}}$. We fix such a pair $(j_{1},j_{2})$. We then factorize the strings $x$, $y$, and $z$ as $x=x_{1}x_{2}x_{3}$, $y=y_{1}y_{2}y_{3}$, and $z=z_{1}z_{2}z_{3}$ so that $s_{j_{1}}$ is the $(x_{1},x_{2})$-boundary, $s_{j_{2}}$ is the $(x_{2},x_{3})$-boundary, $t_{j_{1}}$ is the $(y_{1},y_{2})$-boundary, $t_{j_{2}}$ is the $(y_{2},y_{3})$-boundary, $r_{j_{1}}$ is the $(z_{2},z_{3})$-boundary, and $r_{j_{2}}$ is the $(z_{2},z_{3})$-boundary. It then follows that $(x_{2},y_{2})$ and $(x_{2},z_{2})$ are respectively iterative pairs of $xy$ and $xz$ for $L$. Therefore, the condition (2) follows.
@@
(a) We first consider the case where $s,s^{\prime}]$ is a single turn. We focus on a hill, say, $t_{1},t_{2}]$ that is contained in this turn and claim that this hill has height at most $|Q|^{2}$. This claim is shown as follows. If the hill has height more than $|Q|^{2}$, then there exists an inner state $q\in Q$ that appears in at least $|Q|+1$ state-stack pairs in the left area of the hill. Let $(q,\gamma),(q,\alpha_{1}\gamma),\ldots,(q,\alpha_{[|Q|,1]}\gamma)$ denote those state-stack pairs for appropriately chosen strings $\gamma,\alpha_{1},\ldots,\alpha_{|Q|+1}$. Associated with them, we also choose state-stack pairs $(p_{1},\alpha_{[|Q|,1]}\gamma),\ldots,(p_{|Q|},\alpha_{1}\gamma),(p_{|Q|+1},\gamma)$ in the right area of the hill for certain inner states $p_{1},\ldots,p_{|Q|+1}$. Lemma 5.2(1) then guarantees the existence of a distinct pair $i,j\in[|Q|+1]$ satisfying $p_{i}=p_{j}$. This leads to Case 2, a contradiction.
@@
(3) We estimate the total number of the turns whose bottom heights are a fixed number $h$. Given such a number $h$, we define $bh(h)$ to be the set of all indices $i\in[m-1]$ for which the bottom height of $s_{i},s_{i+1}]$ is exactly $h$. Note that $|bh(0)|\leq|Q|$ follows by (2b). We first assume that $|bh(h+1)|>|\bigcup_{i\in[0,h]_{\mathbb{Z}}}bh(i)|\cdot|Q|$ holds for a certain number $h\in\mathbb{N}$. This implies the existence of more than $|Q|$ turns in $bh(h+1)$, any pair of which contains no turn in $\bigcup_{i\in[0,h]_{\mathbb{Z}}}bh(i)$. Consider $|Q|+1$ state-stack pairs at the boundaries whose heights match the bottom heights of those specific turns. By Lemma 5.3(1), at least two of those state-stack pairs coincide, a contradiction. Hence, we conclude that, for any value $h\in\mathbb{N}$, $|bh(h+1)|\leq|\bigcup_{i\in[0,h]_{\mathbb{Z}}}bh(i)|\cdot|Q|$. We then assert that $|bh(h)|\leq 2^{h-1}|Q|^{h+1}$ for any $h\in\mathbb{N}^{+}$. Since $|bh(i)|\leq 2^{i-1}|Q|^{i+1}$ holds for any $i\in\mathbb{N}^{+}$, by induction hypothesis, it follows that $|bh(h)|\leq|Q|\cdot|\bigcup_{i\in[0,h-1]_{\mathbb{Z}}}bh(i)|\leq|Q|(|Q|+\sum_{i=1}^{h-1}2^{i-1}|Q|^{i+1})\leq(1+\sum_{i=1}^{h-1}2^{i-1})|Q|^{h+1}=2^{h-1}|Q|^{h+1}$.
@@
Basis Case of $d=1$: We begin our proof for the basis case of $d=1$. This comes from Lemma 3.2 by taking an appropriate finite-state 1dpda $M$ for a given $L$ in Lemma 3.1. This completes the proof of the basis case of $d=1$.